# The emergence of galactic thin and thick discs across cosmic history

### v1

@@ 4.3 Adiabatic expansion of the stellar discs
Figure 14: The relative thickness and radial sizes of thin and thick discs in individual two-disc galaxies. The $h_{\mathrm{thin}}/h_{\mathrm{thick}}$ are plotted against $z_{\mathrm{thin}}/z_{\mathrm{thick}}$. The color indicate the mass of the individual galaxies. Galaxies at z=0 and Milky Way are overplotted from Yoachim & Dalcanton (2006); Comerón et al. (2018); Bland-Hawthorn & Gerhard (2016).
@@
The values ($z_{\mathrm{thin}}/z_{\mathrm{thick}}$, $h_{\mathrm{thin}}/h_{\mathrm{thick}}$) are expected to be noisier compared to ($h_{\mathrm{thin}}/z_{\mathrm{thin}}$, $h_{\mathrm{thick}}/z_{\mathrm{thick}}$) due to potential degeneracies and correlations between the denominator and numerator in the fitting process. This correlation may qualitatively support the above disc evolution scenario discussed above e.g. the selective compaction or expansion of thin/thick discs. These processes are likely more efficient than the effects of new gas accretion bringing additional angular momentum and making thin discs longer — a trend that is not observed.
@@ 5 Summary
Despite the dominant sequential picture of thick to thin disc formation revealed in this study, Fig. 11 indicates the co-evolution of the two discs, with the thick disc continually growing as the galaxy grows (although less efficient than thin disc growth). This is in contrast to a simple sequential scenario where two discs form in entirely separate epochs.
We propose that the Toomre-$Q$ self-regulated star formation theory coherently explains the above findings (Sec. 4.1), linking our structural measurements for stellar discs with available gas kinematics of gas disks from recent ALMA and ground-based IFU surveys (Fig. 12). High-mass galaxies achieve lower gas fractions early on, enabling them to host less turbulent gas discs and form thin discs earlier in time. The declining gas fraction over time allows more lower-mass galaxies to form thin discs at later epochs.
@@ Acknowledgements
TT is grateful to the conference organizers and financial support of ELT Science in Light of JWST: The Scientific Landscape for Extremely Large Telescopes in light of JWST held in Miyagi, Japan in June 2024. Attending the conference helped in writing this paper. TT thanks Bruce Elmegreen, Eric Emsellem, Mark Krumholz, Andreas Burkert, Masashi Chiba, Trevor Mendel, Lucas Kimmig, Lucas Valenzuela and Federico Lelli for insightful discussions and Mahsa Kohandel and Hannah Übler for sharing their data (Kohandel et al., 2024; Übler et al., 2019). This work was supported in part by Japan Foundation for Promotion of Astronomy. The data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant No. 140. This research made use of Mathematica.
@@ C.2 Inclination deviation from a perfect edge-on
We quantify the bias introduced by deviations from a perfect edge-on orientation. Slight inclination deviations, of $<7$ deg, expected for our sample (64th percentile) has been shown to minimally affect structural measurements (de Grijs et al., 1997). However, this affect may vary depending on the image quality, the intrinsic properties of the disks, and the presence of a second disc or central bulge. To assess whether inclination effects affect our results we generate simulated images based on the measured structural parameters of our galaxy sample, varying the disc inclination over a range of $\Delta i=[0,25$ deg]. The simulated images were further convolved with the point spread function (PSF) and had Gaussian white noise added from the original variance map. We then refit these convolved simulated images using a model assuming a 90 degree inclination (as adopted in our study) and obtained the fractional biases in the structural parameters, shown in Fig. 19. We use the best fit model classified in Section 3 (e.g., a disk, a disc + bulge, two discs, and two discs + bulge), where single disc, thin disc, and thick disc are denoted in black circle, blue triangle, red square respectively. The error bar shows the standard deviation encompassing our sample of galaxies, representing galaxy to galaxy variation due to the intrinsic structure of disc and image quality for our sample.
As expected the bias is largest for thin disc structures and smallest for thick disc structures. We found 30%, 20% and 10% overestimation are expected for vertical scale height measurements if galaxies have $5$ deg deviation from the perfect edge-on for thin, single, and thick discs respectively. The same value of 30% is found in de Grijs et al. (1997) for a similar intrinsic axial ratio q=0.11 of a thin disk. For galaxies with $q<0.3$ adopted in our sample (see Fig. 3), the assumption of 90 deg lead to median bias of 0.1%, 12%, 1% underestimation for disc scale radii, 17%, 27%, 9% overestimation for disc scale height, and 13%, 27%, 6% underestimation for disc radii/height ratios (values are denoted for single disc, thin disc, and thick disc respectively). These variations contribute to the measured scatter in the reported values. Bizyaev & Mitronova (2009) estimated that the dust extinction correction results in scale height measurements 11% smaller, which may compensate the inclination bias towards smaller values.
@@
Figures 20 and 21 compared the refit results with masking against our fidutial fit results for galaxies classified as single and double disc galaxies, respectively. The comparison shows that the inclusion of masks does not introduce significant systematic biases in the refit results. The scatter observed around the 1:1 line is attributed to the amount of data that were used to constrain the structures. The impact of masking results in a scatter of approximately (9%, 17%, 8%) in $z_{0}$ and (15%, 12%, 15%) in $h_{R}$, which can be considered as an estimate of the statistical and systematic uncertainty. Values in parenthesis are denoted for single disc, thin disc, and thick disc respectively.
Considering the discussions in this section and Appendix C.2, we estimate uncertainties of approximately 30% for $z_{0}$ and 15% $h_{R}$ in our measurements. These uncertainties contribute primarily to the scatter in the correlation, while the median trends are less affected, as the primary sources of uncertainty arises from inclination, and our sample of galaxies statistically peaks at an edge-on orientation due to the strict axial ratio cut as described in Section 2.2.
Figure 20: The measurements for galaxies classified as ‘Single disc’ and ‘disc + bulge’ are compared against fitting results with different masks (see text). For visual aid, the solid line shows the 1:1 relation and the dashed lines show 20% range (of linear quantities) from the 1:1 relation.
Figure 21: The measurements for galaxies classified as ’two discs’ and ’two discs + bugle’ are compared against fitting results with different masks. For visual aid, solid line shows the one-to-one line and the dashed line shows +/- 20% from the one-to-one line in linear scaled measurements.
@@ Appendix D Distribution of galaxy and disc structural parameters
Figure 22: A corner plot of physical properties of host galaxies and individual discs (single, thin, and thick discs).
Figure 23: A corner plot of physical properties of two disc galaxies.
@@ Appendix E M/L variation at observed bands F277W, F356W, and F [^4] W
We compare thin and thick discs mass ratios at redshifts $z=0.1-2$ in observed bands (F227W, F356W, F [^4] W) with a galaxy sample at $z=0$ (Comerón et al., 2014, at Spitzer 3.5 $µm\micron$ band $\approx$ F356W). The derivation of the mass ratio depends on the mass-to-light ratio of thick and thin discs, $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$, for which we assumed a value of 1.2, as derived by Comerón et al. (2011) based on a typical star formation history of the Milky Way, and adopted in Comerón et al. (2014). Figure 24 shows that $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ remains nearly constant across the explored redshift range and exhibits similar values in all three observing bands (F227W, F356W, F [^4] W) with difference being less than 0.25. We compute $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ as a function of redshift at each observing band, with the same star formation histories for thin and thick discs adopted in Comerón et al. (2011). For this computation, we use the python implementation of FSPS (Flexible stellar Population Synthesis) code (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson, 2024). This suggests that it is reasonable to adopt the $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ value at $z=0$ and F356W also for our sample of galaxies at redshifts $z=0.1-2$ in the observed bands.

### v2

@@ 4.3 On why thin discs are smaller than thick discs
Figure 14: The relative thickness and radial sizes of thin and thick discs in individual two-disc galaxies. The ratio $h_{\mathrm{thin}}/h_{\mathrm{thick}}$ is plotted against $z_{\mathrm{thin}}/z_{\mathrm{thick}}$ for each galaxy, represented by black circles filled with color to indicate galaxy mass. Galaxies at z=0 (Yoachim & Dalcanton, 2006; Comerón et al., 2018) and the Milky Way (Bland-Hawthorn & Gerhard, 2016) are overplotted with black dots and a black star, respectively.
@@
The values ($z_{\mathrm{thin}}/z_{\mathrm{thick}}$, $h_{\mathrm{thin}}/h_{\mathrm{thick}}$) are expected to be noisier compared to ($h_{\mathrm{thin}}/z_{\mathrm{thin}}$, $h_{\mathrm{thick}}/z_{\mathrm{thick}}$) due to potential degeneracies and correlations between the denominator and numerator in the fitting process. This correlation may qualitatively support the disc evolution scenario discussed above e.g. the selective compaction or expansion of thin/thick discs. Galaxies may also move within this diagram over time. As thick disc growth slows while the thin disc continues to gain mass and expand in both height and radius (inside-out disc evolution; Minchev et al., 2013), data points may shift from the lower left to the upper right along the sequence. In the downsizing thin disc formation, more massive galaxies might have had more time for their thin discs to grow; however, such trends are not clearly evident in the diagram. Constraining stellar ages could further help test these scenarios by revealing how galaxies evolve within the diagram.
The formation of a smaller thin disc within an existing thick disc suggests an outside-in transformation when the thin disc emerges in pre-existing thick disc. Tadaki et al. 2020 found that in massive star-forming galaxies at $z\sim 2$ ($M_{*}>10^{11}M_{\odot}$), the star-forming region is about half the size of the pre-existing stellar disc, possibly marking the onset of thin disc formation. According to the downsizing scenario, such massive galaxies may have formed thin discs earlier than the most massive galaxies in our sample ($\sim 8$ Gyr ago; see Fig. 9).
4.4 Origin of thick and thin discs
In this section, we discuss which formation scenarios of thin and thick discs are supported by our findings. By directly observing galaxies in the past, we confirm sequential stellar disc formation: galaxies first form a thick disc and later evolve into two-disc composite galaxies by forming a subsequent thin disc from within. The accretion of ex situ stars from small satellites cannot be the primary mechanism for the thick disc formation. Bringing satellites to the galactic disc scale via dynamical friction, in the absence of a pre-existing disc, would take longer (Peñarrubia et al., 2002; Villalobos & Helmi, 2008) than the observed early appearance of dissipative gas discs at $\sim$4-7 (as early as $700$ Myr, Neeleman et al. 2020; Rizzo et al. 2020; Lelli et al. 2021; Tsukui & Iguchi 2021; Rowland et al. 2024). These early discs likely form through gas-rich mergers and/or cold gas accretion. Similarly, slow "progressive thickening" is inconsistent with our observations, as galaxies seem to have thick discs as early as $z\sim 3$. However, it may still be viable if the process occurs on a much shorter timescale than our observations can capture.
Over the long term, these mechanisms may contribute to the growth of the thick disc. Both thin and thick disc masses increase with galaxy mass, following different power-low slopes (Fig.11). If we interpret this as galaxies evolving along the tracks, they continue increasing their thick disc mass (albeit less efficiently than their thin discs) after forming a thin disc. Satellite accretion and heating may be a viable mechanism for thick disc growth at later stages (Pinna et al., 2019a, b; Martig et al., 2021), driving the continuous growth of thick discs as total galaxy mass increases.
We demonstrate that sequential formation, thick then thin, proceeds in a downsizing manner, where more massive galaxies form thin discs earlier. By linking to high-$z$ gas kinematics, we highlight the role of ISM turbulence in determining the timing of thin disc formation (Sec. 4.1). This is consistent with the scenario in which thick discs rapidly form in chaotic gas-rich turbulent disc and thin discs subsequently form from quiescent low gas-fraction discs (Beraldo e Silva et al., 2021; Yu et al., 2023; Bland-Hawthorn et al., 2025) [^10] In gas-rich turbulent discs, Yu et al. (2023) suggest that thick disc stars form in hot orbits (”born-hot”), whereas Beraldo e Silva et al. (2021); Bland-Hawthorn et al. (2025) propose that most stars form near the disc mid-plane and are quickly heated (clumps and sloshing, ”instant thickening”). The structural analysis in this paper cannot distinguish them, but color gradients or rest-frame ultra violet (UV) observations for a single (thick) disc in high-z galaxies may help differentiate them.. Under gravitational stability-regulated disc formation, the gas turbulence is related to gas fraction $f_{\mathrm{gas}}$ by $v/\sigma\propto 1/f_{\mathrm{gas}}$ (Genzel et al., 2011), suggesting that higher $f_{\mathrm{gas}}$ leads to higher turbulence (lower $v/\sigma$). High turbulence pressure may prohibit the thin disc formation (van Donkelaar et al., 2022). In cosmic downsizing, more massive galaxies convert gas into stars more efficiently (Behroozi et al., 2013), forming thick discs earlier (Comerón, 2021). As they achieve lower $f_{\mathrm{gas}}$ and lower turbulence, they transition to form thin discs earlier.
Archaeological studies of nearby edge-on galaxies also support the downsizing picture. More massive galaxies tend to have older thick discs with higher [$\alpha$/Fe] than low-mass galaxies, indicating rapid and early thick disc formation in massive galaxies (Sattler et al., 2024, 2023; Pinna et al., 2019a, b; Martig et al., 2021). Ground-based observations support this view showing that disc turbulence and $f_{\mathrm{gas}}$ increases at higher redshifts (Förster Schreiber et al., 2009; Genzel et al., 2011; Wisnioski et al., 2015; Übler et al., 2019; Tacconi et al., 2020; Rizzo et al., 2024). Although recent ALMA observations have revealed surprisingly low relative turbulence, with $v/\sigma$ values as high as $\sim 10$ at $\sim$4, this remain consistent with this downsizing framework (Rizzo et al., 2020; Lelli et al., 2021). The available spatially resolved kinematic measurements are generally biased towards massive systems (Fig. 12), which may already host substantial thick discs and thus exhibit lower $f_{\mathrm{gas}}$. Interestingly, some ALMA detected lower mass galaxies (e.g., Neeleman et al., 2020; Tsukui & Iguchi, 2021; Parlanti et al., 2023) show enhanced turbulence (low $v/\sigma$) and high gas fractions, suggesting ongoing thick disc formation. For example, BRI 1335-0417 has a high gas fraction of $\sim$70and turbulence $v/\sigma\sim 2.5\pm 0.5$, corresponding to an axial ratio $q\sim 5$, clearly placing it in the thick disc formation regime (Fig. 12). Additionally, it uniquely shows spiral and bar structures (Tsukui & Iguchi, 2021; Tsukui et al., 2024).
Understanding the role of gas-rich turbulent discs in thick disc formation - and the puzzling presence of spiral and bar in such environments - has advanced significantly through numerical simulations. Recent simulations of gas-dominated discs show that high gas-fraction discs can rapidly develop spirals and bars, while young stellar bars form through disc shear flows (Bland-Hawthorn et al., 2024). These simulations also reveal an intriguing mechanism: stochastic star formation within complex gas substructures induces bulk motion (sloshing) of the gas disc relative to the halo potential, dispersing stars. The energy from this bulk motion is transferred to the stars, contributing to thick disc formation (Bland-Hawthorn et al., 2025). In these early epochs, clumps may have also contributed to stellar scattering (Beraldo e Silva et al., 2020, 2021).
A potential contradiction to our findings "thick disc first, thin disc later" is the presence of an old, metal-poor thin disc in the Milky Way (Nepal et al., 2024). However, this low-mass component may have formed through later satellite accretion events, where dynamical friction drags satellites into the disc plane, preferentially from prograde satellites (Walker et al., 1996) , or it may consist of stars that survived heating in the gas-rich disc (Beraldo e Silva et al., 2021; Yu et al., 2023; Bland-Hawthorn et al., 2025).
The relative importance of different growth mechanisms likely depends on a galaxy’s properties (e.g., total mass at a given epoch) and formation history (e.g., Pinna et al., 2024; Yu et al., 2021). This study demonstrates the JWST potential to understand the Milky Way’s formation history by directly examining Milky Way-sized progenitors at earlier epochs (Fig. 9) and determine if the Milky Way has a distinct formation history compared to others (e.g., Rey et al., 2023).
@@ 5 Summary
Despite the dominant sequential picture of thick to thin disc formation revealed in this study, Fig. 11 indicates the co-evolution of the two discs, with the thick disc continuously growing as the galaxy grows (although less efficient than thin disc growth). This is in contrast to a simple sequential scenario where two discs form in entirely separate epochs.
We propose that the Toomre-$Q$ self-regulated star formation coherently explains the above findings (Sec. 4.1), linking our structural measurements for stellar discs with available gas kinematics of gas discs from recent ALMA and ground-based IFU surveys (Fig. 12). High-mass galaxies achieve lower gas fractions early on, enabling them to host less turbulent gas discs and form thin discs earlier in time. The declining gas fraction over time allows more lower-mass galaxies to form thin discs at later epochs.
@@ Acknowledgements
We are grateful to the anonymous reviewer for their constructive feedback, which has significantly improved the clarity and quality of this paper. TT is grateful to the conference organizers and financial support of the ELT Science in Light of JWST meeting, held in Miyagi, Japan in June 2024. Attending the conference helped to focus this research and the eventual paper. TT thanks Bruce Elmegreen, Eric Emsellem, Mark Krumholz, Andreas Burkert, Masashi Chiba, Trevor Mendel, Lucas Kimmig, Lucas Valenzuela and Federico Lelli for insightful discussions and Mahsa Kohandel and Hannah Übler for sharing their data (Kohandel et al., 2024; Übler et al., 2019). This research was supported by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013. This work was supported in part by Japan Foundation for Promotion of Astronomy. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program # 1180, 1210, 1895, 2079, 2514, 1181, 3577, 1345, 2750, 2279, 1727, 1837, 1283, 2198. The data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant No. 140. This work made use of the following software packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez & Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (Wes McKinney, 2010; pandas development team, 2024), python (Van Rossum & Drake, 2009), scipy (Virtanen et al., 2020; Gommers et al., 2025), Cython, and scikit-image (van der Walt et al., 2014). This research has made use of NASA’s Astrophysics Data System. This research made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2024) and Mathematica (Wolfram Research, 2024). Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2024).
@@ C.2 Inclination deviation from a perfect edge-on
We quantify the bias introduced by deviations from a perfect edge-on orientation. Slight inclination deviations, of $<7$ deg, expected for our sample (64th percentile) has been shown to minimally affect structural measurements (de Grijs et al., 1997). However, this affect may vary depending on the image quality, the intrinsic properties of the discs, and the presence of a second disc or central bulge. To assess whether inclination effects affect our results we generate simulated images based on the measured structural parameters of our galaxy sample (111 best-fit models), varying the disc inclination over a range of $\Delta i=[0,15$ deg] in 0.5 deg intervals. The simulated images were further convolved with the PSF and had Gaussian white noise added from the original variance map. We then refit these convolved simulated images using a model assuming a 90 degree inclination (as adopted in our study) and obtained the fractional biases in the structural parameters, shown in Fig. 19. We use the best fit model classified in Section 3 (e.g., a disc, a disc + bulge, two discs, and two discs + bulge), where single disc, thin disc, and thick disc are denoted in black circle, blue triangle, red square respectively. The error bar shows the standard deviation encompassing our sample of galaxies, representing galaxy to galaxy variation due to the intrinsic structure of disc and image quality for our sample.
As expected the bias is largest for thin disc structures and smallest for thick disc structures. We found 30%, 20% and 10% overestimation are expected for vertical scale height measurements if galaxies have $5$ deg deviation from the perfect edge-on for thin, single, and thick discs respectively. The same value of 30% is found in de Grijs et al. (1997) for a similar intrinsic axial ratio q=0.11 of a thin disc. For galaxies with $q<0.3$ adopted in our sample (see Fig. 3), the assumption of 90 deg lead to median bias of 0.1%, 12%, 1% underestimation for disc scale radii, 17%, 27%, 9% overestimation for disc scale height, and 13%, 27%, 6% underestimation for disc radii/height ratios (values are denoted for single disc, thin disc, and thick disc respectively). These variations contribute to the measured scatter in the reported values.
@@
Figures 20 and 21 compare the refit results with masking against our fiducial fit results for galaxies classified as single- and double-disc galaxies, respectively. The impact of masking results is summarized by the table 3. The scatter and mean values quantify the deviation introduced by masking, relative to the fiducial (unmasked) fit. The comparison shows that masking does not introduce a significant systematic shift in the refit results, as the mean deviation is smaller than the introduced scatter. This suggests that the reduced amount of available data to constrain the structural parameters has a greater effect than the structural differences in the masked regions. As expected, the midplane mask primarily affects disc heights, while the truncation mask has a greater impact on disc radii. This validates the fiducial fit and confirms that the potential midplane disc structure and disc truncations do not affect our measurements significantly. The moderate mean deviation for height measurements using midplane mask is consistent with the dust effect $\sim 11\%$ estimated in similar bands (Bizyaev & Mitronova, 2009).
Table 3: Fractional scatter and mean deviation of measurements obtained using different masks relative to the fiducial (unmasked) fit.
Parameter Disc Component Mask 1 (Bulge) Mask 2 (Bulge + Midplane) Mask 3 (Bulge + Truncation)
Scatter / Mean dev. (%) Scatter / Mean dev. (%) Scatter / Mean dev. (%)
$z_{\text{single}}$ Single Disc 8.21 / -2.59 8.78 / -6.04 8.82 / -3.09
$z_{\text{thin}}$ Thin Disc 11.07 / -1.30 22.29 / -14.45 19.34 / 0.02
$z_{\text{thick}}$ Thick Disc 6.79 / -0.94 8.90 / -3.40 7.37 / 1.34
$h_{\text{single}}$ Single Disc 12.16 / 1.62 12.47 / -0.49 21.92 / -6.70
$h_{\text{thin}}$ Thin Disc 7.62 / 0.47 11.61 / -1.59 16.69 / 1.96
$h_{\text{thick}}$ Thick Disc 3.01 / 0.79 4.54 / -2.89 40.26 / -14.31
Considering the discussions in this section and Appendix C.2, the systematic uncertainty from the disc inclination dominates in our measurements. The thin disc is most affected by the inclination effect, introducing a median bias of approximately 12% underestimation and 27% overestimation in the measured disc size and height, with smaller biases for single and thick discs. Therefore, the systematic uncertainty is minor compared to the dynamic range of the measurements and seen trend (e.g., Fig. 5 and 6). The visual exclusion of face-on discs with dust morphology further reduce the uncertainty than this estimate. These uncertainties contribute primarily to the scatter in the correlation, while the median trends are less affected.
Figure 20: The measurements for galaxies classified as ‘Single disc’ and ‘disc + bulge’ are compared against fitting results with different masks: mask1 (‘bulge mask’), mask2 (‘bulge mask’ + ‘midplane mask’) and, mask3 (‘bulge mask’ + ‘disc truncation mask’). See the text for details. For visual aid, the solid line shows the 1:1 relation and the dashed lines show 20% range (of linear quantities) from the 1:1 relation.
Figure 21: The measurements for galaxies classified as ’two discs’ and ’two discs + bugle’ are compared against fitting results with different masks (same as Fig. 20). For visual aid, solid line shows the one-to-one line and the dashed line shows +/- 20% from the one-to-one line in linear scaled measurements.
@@ Appendix D Distribution of galaxy and disc structural parameters
Figure 22: A corner plot of physical properties of host galaxies and individual discs, with single, thin, and thick discs represented by black, blue, and red points, respectively.
Figure 23: A corner plot of physical properties of two disc galaxies (blue-red points).
@@ Appendix E M/L variation at observed bands F277W, F356W, and F [^4] W
In Sec. 3.4, We compare thin and thick discs mass ratios at redshifts $z=0.1-2$ in observed bands (F227W, F356W, F [^4] W) with a galaxy sample at $z=0$ (Comerón et al., 2014, at Spitzer 3.5 $µm\micron$ band $\approx$ F356W). The derivation of the mass ratio depends on the mass-to-light ratio of thick and thin discs, $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$, for which we assumed a value of 1.2, as derived by Comerón et al. (2011) based on a typical star formation history of the Milky Way, and adopted in Comerón et al. (2014).
A key question is whether this value can be used for galaxies with different redshift and observed at different bands although close to the band used in (Comerón et al., 2011). To address this, we compute $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ as a function of redshift at each observing band, with the same four star formation histories for thin and thick discs adopted in Comerón et al. (2011). For this computation, we use the python implementation of FSPS (Flexible stellar Population Synthesis) code (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson, 2024). Figure 24 shows that all $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ remains nearly constant across the explored redshift range and exhibits similar values in all three observing bands (F227W, F356W, F [^4] W) with difference being less than 0.25. This suggests that it is reasonable to adopt the $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}$ value at $z=0$ and F356W also for our sample of galaxies at redshifts $z=0.1-2$ in the observed bands.